OpenStudy (anonymous):

the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

Use: \[\large y-y_1 = m(x-x_1)\] here m = 2 x1 = -1 and y1 = 5..

OpenStudy (anonymous):

Plug in the values and then solve: \[y - 5 = 2(x + 1) \implies y - 5 = x + 2 \implies x - y = -7\]